OpenStudy (anonymous):

Write an equation of the line containing the given point and parallel to the given line. (-9,2); 8x=7y+5

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

1. get the equation into y=mx+b form. y= 8/7 x-5/7 2. parallel means same slope, so your new equation will be y=8/7 x+b 3. To find the b, plug in your x and your y and solve for b 2=8/7(-9)+b 2=-72/7 +b 2+72/7=b 86/7=b 4. Now rewrite your equation. y= 8/7 x+ 86/7 weird looking, but I'm pretty sure it's right. (If you divide out the sevens you get a whole ton of crazy decimals, so it's easier to leave it in fractional form).